PICOSECOND TIME-RESOLVED IR-UV PUMP-PROBE SPECTROSCOPIC STUDY ON VIBRATIONAL ENERGY RELAXATION OF BENZENE DIMER AND TRIMER IN THE CH STRETCHING REGION

\underline{RYOJI KUSAKA}, YOSHIYA INOKUCHI, and TAKAYUKI EBATA, Department of Chemistry, Graduate School of Science, Hiroshima University, Higashi-Hiroshima, 739-8526, Japan.

Vibrational energy relaxation (VER) in the CH stretching region of benzene dimer (Bz_{2}) and trimer (Bz_{3}) has been studied by IR-UV pump-probe spectroscopy in supersonic beams. Firstly, we investigated isotope-substituted hd heterodimer, where h=C_{6}H_{6} and d=C_{6}D_{6}, because the Stem and Top sites in the hd dimer can be site-selectively excited, different from hh homodimer. The two h(stem)d(top) and h(top)d(stem) isomers show remarkable difference in the lifetimes of intracluster vibrational energy redistribution (IVR). In the transient UV spectra, we observed a broad electronic transition due to the bath modes. The time evolutions of the bath modes can be described by a three step VER model involving IVR and vibrational predissociation (VP). This model was also confirmed by the observed rise profile of the Bz fragment. Secondly, we investigated hh homodimer. The hh homodimer shows the stepwise VER process with time constants similar to those of the hd dimer, suggesting a very weak excitation-exchange coupling of the vibrations between the two sites of the hh dimer. Finally, we found that Bz_{3} also exhibits the stepwise VER process, though each step is faster than Bz_{2}.